If I am solving for "Y" in this equation, how would I do that?
6x-3y=12
Solve for y
when x=2y

Answer by checkley77(12844)

6x-3y=12
Solve for y
when x=2y
6(2Y)-3Y=12
12Y-3Y=12
9Y=12
Y=12/9
Y=4/3 ANS.
6(2*4/3)-3*4/3=12
6(8/3)-12/3=12
2*8-4=12
16-4=12
12=12